### v1

# In-Context Multi-Operator Learning with DeepOSets

Abstract
In-context Learning (ICL) is the remarkable capability displayed by some machine learning models to learn from examples in a prompt, without any further weight updates. ICL had originally been thought to emerge from the self-attention mechanism in autoregressive transformer architectures. DeepOSets is a non-autoregressive, non-attention based neural architecture that combines set learning via the DeepSets architecture with operator learning via Deep Operator Networks (DeepONets). In a previous study, DeepOSets was shown to display ICL capabilities in supervised learning problems. In this paper, we show that the DeepOSets architecture, with the appropriate modifications, is a multi-operator in-context learner that can recover the solution operator of a new PDE, not seen during training, from example pairs of parameter and solution placed in a user prompt, without any weight updates. Furthermore, we show that DeepOSets is a universal uniform approximator over a class of continuous operators, which we believe is the first result of its kind in the literature of scientific machine learning. This means that a single DeepOSets architecture exists that approximates in-context any continuous operator in the class to any fixed desired degree accuracy, given an appropriate number of examples in the prompt. Experiments with Poisson and reaction-diffusion forward and inverse boundary-value problems demonstrate the ability of the proposed model to use in-context examples to predict accurately the solutions corresponding to parameter queries for PDEs not seen during training.
1 Introduction
In learning the solution operator of a parametric PDE from sample data, one is not interested in merely obtaining a single solution for a single PDE, but to learn the mapping between parameter and solution, so that new solutions corresponding to new values of the parameter can be obtained quickly and without further training. Furthermore, both the parameter and the solution are typically functions in infinite-dimensional spaces. For example, the problem might be to learn the space-varying fluid pressure (solution) corresponding to the space-varying permeability in a porous medium (parameter), where the physical process is governed by Darcy’s Equation. The training data in this case consists of pairs of parameter and corresponding solution, obtained from experiments or, more commonly, by solving Darcy’s equation using a high-fidelity solver. The trained operator allows quick exploration of the parameter space, which is unfeasible to do with expensive experiments or high-fidelity solves. The most well-known operator learners are DeepONets ludeeponet2021 and Fourier Neural Operators li2020fourier, which are both based on specific deep neural network architectures.
A more recent development is foundation models for scientific problems, which is inspired by the success of of large language models. A foundation model can be trained on a large body of data and be adapted for new tasks, perhaps with no further training bommasani2021opportunities. In the context of PDEs, foundation models correspond to multiple operator (multi-operator) learning. The challenge could be to be able to train a model with data from diverse PDEs with diverse equation coefficients and boundary conditions, and then produce fast solutions for new PDEs on demand. In contrast to single operator learning, here a well-posedness question arises, since multiple solution operators acting on the same parameter function will produce different solutions. How is the model supposed to identify the correct operator? There are two basic methods to resolve this issue: the first, is to fine-tune the model on a small amount of data from the new PDE; the second is to use in-context learning (ICL), when the PDE is identified from example pairs in a prompt with no further training.
In-Context Operator Networks (ICON) yang2023context and its subsequent variants yang2024pde; yang2023ICON-LM; zhang2025GenICON remain the only published approach to in-context learning of PDE solution operators (see Section 2 for other work related to multi-operator learning and foundation models for PDEs). ICON is a framework that adapts GPT-style in-context learning to PDE problems, by employing a decoder transformer to generate the solution from a sequential prompt followed by the query. ICON is thus an autoregressive, attention-based approach, with quadratic complexity in prompt length, due to the dot-product self-attention mechanism. This can introduce significant computational overhead and loss of accuracy with very long prompts. In addition, if the query changes, the prompt has to be processed sequentially again from the beginning (unless caching is used).
In this paper, we propose an alternative approach to in-context learning of PDE solution operators, which processes the prompt non-autoregressively, in a fully parallel manner, without using self-attention, being therefore much faster during both training and inference time. This approach is based on DeepOSets, a neural architecture that combines set learning via the DeepSets architecture zaheer2017deep with operator learning via the previously mentioned DeepONets architecture. In DeepOSets, the DeepSets layer acts as an encoder of the prompt into a latent representation that is the input, along with the query, to the branch path of a DeepONet. The DeepSets layer confers linear complexity in the length of the prompt, in contrast to the quadratic complexity of self-attention, and a built-in permutation-invariance inductive bias to the order of the examples in the prompt.
DeepOSets was shown in a previous publication to display ICL capabilities in regression problems chiu2024deeposets. Here, we introduce architectural modifications to DeepOSets that allow it to act as a multi-operator learner for parametric PDEs. This is accomplished by discretizing a variable number of solution and parameter function pairs, which are entered into the prompt and encoded by the DeepSets layer into latent representation. The latter is concatenated to the discretization of the query parameter function and entered directly into the branch path of the DeepONet.
Crucially, we prove that this DeepOSets architecture approximates uniformly a compact set of continuous solution operators, which means that DeepOSets are not only universal approximators of single solution operators, like DeepONets, but they are universal approximators of multiple operators, in the sense that a single DeepOSets architecture exists that can approximate any operator in the class to the same fixed degree of accuracy. This guarantees that DeepOSets can learn in-context the solution operators of multiple PDEs.
Experiments with Poisson and reaction-diffusion forward and inverse problems demonstrate that DeepOSets can accurately predict the solution to a new PDE with inverse or forward problem and unknown value of equation coefficients and boundary conditions learned in-context.
2 Related Work
MIONet is probably the first approach to attempt multi-input operator learning, by encoding several input functions via multiple branch network (and a shared trunk network) of a DeepONet, though it still learns a single operator jin2022mionet. POSEIDON is a large foundation model for PDEs based on a scalable Operator Transformer (scOT) with a vision transformer architecture, is pretrained on compressible Euler and incompressible Navier-Stokes equations and applied to 15 downstream tasks, showing better sample efficiency and accuracy than baselines herde2024poseidon. PROSE-PDE is a transformer-based multi-operator learning approach trained on one-dimensional time-dependent nonlinear PDEs that encodes both numerical and symbolic data types for forward and inverse PDE problems sun2025towards. LeMON is based on pretraining and fine-tuning for multi-operator learning, and introduces low-rank adaptation to reduce computational cost sun2024lemon. MT-DeepONet introduces modifications to the branch network of a DeepONet to accommodate various functional forms of parameterized coefficients and multiple geometries kumar2025synergistic. In-context learning for PDE solution operators was introduced by ICON yang2023context. The follow-up study on 1D scalar conservation laws provides a detailed methodology for applying ICON to temporally evolving PDEs yang2024pde, while ICON-LM integrates ICON concepts with language-model architectures yang2023ICON-LM. GenICON introduces a generative, probabilistic approach to ICON zhang2025GenICON.
3 The DeepOSets Architecture for In-Context Operator Learning
The DeepOSets model is a hybrid neural network comprising a DeepSets encoder and a DeepONet decoder, as seen in Figure 1. The DeepSets module encodes non-autoregressively the discretized parameter and solution function example pairs in the prompt. Each pair $(u_{i},G(u_{i}))$ is first mapped to an individual fixed-length embedding $\bf{h}_{i}$ by a shared multi-layer perceptron (MLP) denoted $\phi$. These intermediate representations are then combined by a pooling layer into the embedding vector $\bf{h}$. This introduces a permutation invariance inductive bias to the prompt examples, as long as the pooling layer is a permutation invariant operation, such as min, max, mean or set attention pooling. The DeepONet takes the concatenation of the embedding $\bf{h}$ and the discretized query parameter function $u_{q}$ at its branch path and the query coordinates $x_{q}$ at its trunk path, and outputs the predicted solution function $\hat{G}(u_{q})$ evaluated at $x_{q}$. The output is therefore entirely mesh-free.
Figure 1: DeepOSets architecture for in-context multi-operator learning.
4 Universality of DeepOSets for In-Context Operator Learning
In this section, we prove that DeepOSets are universal uniform approximators over a compact set of continuous operators, i.e., a single DeepOSets neural network exists that bounds the supremum of the error over the entire class to any desired fixed accuracy. This shows that a unique DeepOSets neural network could be learned that approximates in-context any continuous operator to the desired accuracy, given an appropriate number of examples in the prompt.
We first define the class of operators of interest. Let $X:=C(K_{1}),~Y:=C(K_{2})$, with $K_{1},K_{2}\subset\mathbb{R}^{d}$ compact subsets of some Euclidean space. Let $V\subset X$ be compact. Given this, consider the space of continuous operators (in the induced operator norm):
$\mathcal{O}:=\{G:V\rightarrow Y~\text{continuous}\}$
with norm given by
$\|G\|_{\mathcal{O}}:=\sup_{x\in V}{\|G(x)\|_{Y}}.$
We let $\mathcal{K}\subset\mathcal{O}$ to be compact subset of this space of operators. We will need to discretize functions in $X$ on different grids in $K_{1}$ and will also need to consider certain coverings of the set $V$. For this purpose, we introduce the following definition.
Definition 4.1 (Grid).
Let $\delta>0$ and $C>1$. A $(\delta,C)$-discretization of a compact metric space $K$ is a set of points $\textbf{x}=\{x_{1},\cdots,x_{n}\}$ such that
$|\{x_{i}\in B(x,\delta/2)\}|<C|\{x_{i}\in B(y,\delta)\}|$
for any $x,y\in K$.
Intuitively, a $(\delta,C)$-discretization is a $\delta$-net of $K$ which is not too unbalanced, in the sense that the density of points is relatively uniform. Let us first note that this definition is not vacuous.
Lemma 4.1.
Let $K$ be a compact set in a metric space, $\delta>0$ and $C>1$. Then there always exists a $(\delta,C)$-discretization of $K$.
Proof.
See the Appendix.∎
Given a continuous function $u:K\rightarrow\mathbb{R}$ and a $(\delta,C)$-discretization $\textbf{x}=(x_{1},...,x_{n})$, we denote by
$u(\textbf{x}):=\left(u(x_{i})\right)_{i=1}^{n}\in\mathbb{R}^{n}$ (1)
for the function evaluated at the grid x. Given these definitions and notation, our main result is the following.
Theorem 4.2 (Uniform universality for In-Context Operator Learning).
Given $\varepsilon>0$, $C>1$ and a compact set $\mathcal{K}\subset\mathcal{O}$, there exist integers $N,M\geq 1$, $\delta>0$, fixed discretizations $\textbf{x}=\{x_{1},...,x_{k}\}$ of $K_{1}$ and $\{y_{1},...,y_{n}\}$ of $K_{2}$, and continuous functions $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{M}$, $\rho:\mathbb{R}^{k+N}\rightarrow\mathbb{R}^{M}$, and $\Phi:\mathbb{R}^{k+n}\rightarrow\mathbb{R}^{N}$, such that
$\sup_{G\in\mathcal{K}}\>\left|\,G(u_{q})(y)-\Psi(y)\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\right|<\varepsilon.$
for any $(\delta,C)$-discretization $\textbf{u}:=(u_{1},\cdots,u_{m})$ of $V$, any $u_{q}\in V$, and any $y\in K_{2}$.
Proof.
See the Appendix.∎
The continuous functions guaranteed by Theorem 4.2 can be approximated using neural networks by invoking standard universal approximation theorems for MLPs cybenko1989approximation; hornik1991approximation; pinkus1999approximation. Here it is important that we are averaging the outputs of $\Phi$ to ensure that the function $\rho$ only needs to be approximated on a compact set.
5 Experimental Results
Here, we demonstrate that the proposed DeepOSets architecture can learn the solution operator for multiple PDEs, with the correct PDE being disambiguated with the prompt examples, without any further weight updates. We consider two one-dimensional parametric PDEs and forward and inverse problems for each.
The Poisson boundary-value PDE reads:
$\frac{d^{2}u(x)}{dx^{2}}$ $=\,f(x)\,,\quad 0<x<1\,$, (2)
with $u(0)=u_{0}$ and $u(1)=u_{1}$. The forward problem is to find the solution $u(x)$ given the forcing function $f(x)$. The inverse or control problem is to find the forcing function $f(x)$ that produces the given trajectory $u(x)$. Very importantly, the boundary values $u_{0}$ and $u_{1}$ are unknown and must be learned in-context.
The reaction-diffusion boundary-value PDE is:
$a\frac{d^{2}u(x)}{dx^{2}}+k(x)u(x)=c\,,\quad 0<x<1\,$, (3)
with $u(0)=u_{0}$ and $u(1)=u_{1}$. The forward problem is to find the solution $u(x)$ given the coefficient function $k(x)$. The inverse problem is to find the coefficient function $k(x)$ that produces the given solution $u(x)$. Here, the coefficients $a$ and $c$ and the boundary values $u_{0}$ and $u_{1}$ are all unknown and must be learned in-context.
(a)         Forward Poisson
    $\frac{d^{2}{\color[rgb]{1,0,0}G(u)(x)}}{dx^{2}}={\color[rgb]{0,0,1}u(x)}$, given $u(0),u(1)$
(b)         Inverse Poisson
    $\frac{d^{2}{\color[rgb]{0,0,1}u(x)}}{dx^{2}}={\color[rgb]{1,0,0}G(u)(x)}$
, given $u(0),u(1)$
(c)      Forward Reaction-Diffusion
   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{1,0,0}u(x)}+{\color[rgb]{0,0,1}k(x)}{\color[rgb]{1,0,0}u(x)}=c$, given $a$, $c$
(d)      Inverse Reaction-Diffusion
   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{0,0,1}u(x)}+{\color[rgb]{1,0,0}k(x)}{\color[rgb]{0,0,1}u(x)}=c$, given $a$, $c$
Figure 2: In-context PDE solution operator learning with the DeepOSets architecture. In each case, the top plot contains the parameter functions while the bottom plot contains the corresponding solution functions. The gray lines indicate in each case the parameter-solution function pairs given in the prompt. The blue line indicates the query parameter function, while the red line indicates the corresponding solution predicted by the trained DeepOSets model. The black line indicates the exact solution corresponding to the query, using the appropriate PDE. The DeepOsets model is trained to learn the four PDE problems simultaneously, with coefficients sampled randomly.
All PDE coefficients and boundary values are chosen randomly, so that the number of solution operators involved is in fact infinite. The physical domain in all cases is the interval $[0,1]$, which is discretized uniformly into 100 points. The prompts during both training and testing contain 4 example pairs of parameter and solution functions. The MLP $\phi$ in the DeepOSets contains 2 hidden layers of 400 neurons each, and the latent vector encoding the prompt data has size 500. The pooling operation is the sample mean. The branch and trunk networks in the DeepOSets contain hidden layers of 200, 400, and 200 neurons with intermediate layer of 150 neurons. All neural network use ReLU as the activation function.
To generate the training data, for each of the four problems (Poisson/reaction-fiffusion, forward/inverse), 100 parameter functions are sampled independently from a Gaussian process with the squared-exponential kernel (discretized on the uniform 100-point grid) with variance 2 and length scale $0.5$. For each parameter function and PDE, the boundary values $u_{0}$ and $u_{1}$ are sampled uniformly in the interval [1,1][-1,1]. In addition, in the case of the reaction-diffusion PDE, the coefficients $a$ and $c$ are sampled uniformly in the intervals $[0.025,0.075]$ and [2,2][-2,2], respectively. Then the PDE is solved for the given parameter function, boundary values, and coefficients to obtain the groundtruth solutions. The model was trained by minimizing the mean squared error loss between predicted and groundtruth solutions over a mini-batch of 100 operators sampled randomly from the training data at each iteration, using the Adam optimizer with learning rate 1e-3, for 50k iterations. Training took 14min, achieving a training loss 8.4e-3.
Sample results are shown in Figure 2. At inference time, DeepOSets can perform accurate in-context learning, without knowing which PDE, inverse or forward problem, or the value of the equation coefficients and boundary conditions are. All these unknowns are correctly disambiguated from the data in the prompt, and the solution corresponding to the query parameter is accurately predicted, without any weight updates.
6 Conclusion
We have presented DeepOSets for in-context multi-operator learning, a novel non-autoregressive, non-attention-based approach to learning solution operators of multiple parametric PDEs. By combining DeepSets for set learning with DeepONets for operator learning, DeepOSets addresses key limitations of existing approaches while maintaining strong in-context learning capabilities. The proposed architecture offers several advantages: fully parallel processing of prompts enables faster training and inference compared to sequential methods; linear complexity in prompt length provides better scalability than quadratic self-attention mechanisms; and built-in permutation-invariance provides beneficial inductive bias for training efficiency. We prove that DeepOSets are universal uniform approximators over a class of operators, which we believe is the first result of its kind to be published in the Scientific Machine Learning area. Our experiments on Poisson and reaction-diffusion problems demonstrate that DeepOSets successfully disambiguates between multiple PDEs, forward and inverse problems, and unknown equation coefficients and boundary conditions purely from prompted examples. The model achieves accurate predictions without weight updates at inference time, requiring only 14 minutes of training on all four problems simultaneously. Future work will extend DeepOSets to higher-dimensional PDEs, more diverse equation families, and longer prompts. DeepOSets establishes a computationally efficient foundation for building general-purpose PDE solvers capable of in-context adaptation, contributing toward foundation models for scientific computing.
Appendix
Let us begin by proving Lemma 4.1 showing that $(\delta,C)$-discretizations always exist.
Proof of Lemma 4.1.
Let $K\subset X$ be compact, and let $\delta>0$ be given. Choose a maximal $\delta$-separated subset of $K$, i.e., a subset $\mathcal{X}_{\delta}$ such that $d(x_{i},x_{j})\geq\delta$ for all $i\neq j$, and which is maximal with respect to this property. It is clear that the balls of radius $\delta$ around the points $x_{i}$ give a minimal covering of $K$, and so by compactness $\mathcal{X}_{\delta}$ must be a finite set. We claim that $\mathcal{X}_{\delta}$ is a $(\delta,C)$-covering for any $C>1$.
First, note that any ball of radius $\delta/2$ contains at most one point of $\mathcal{X}_{\delta}$. Indeed, fix $z\in K$ and suppose $x_{i},x_{j}\in\mathcal{X}_{\delta}$ both lie in $B(z,\delta/2)$ with $i\neq j$. Then, by the triangle inequality,
$d(x_{i},x_{j})\;\leq\;d(x_{i},z)+d(z,x_{j})\;<\;\frac{\delta}{2}+\frac{\delta}{2}\;=\;\delta,$
which contradicts the $\delta$–separation of $\mathcal{X}_{\delta}$. Hence, for every $z\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(z,\delta/2)\}\bigr|\leq 1.$
Next, since balls of radius $\delta$ around the points in $\mathcal{X}_{\delta}$ cover $K$, for every $y\in K$ there exists some $x_{j}\in\mathcal{X}_{\delta}$ with $d(y,x_{j})<\delta$, i.e. $x_{j}\in B(y,\delta)$. Thus, for all $y\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|\geq 1.$
Combining these two observations, we obtain, for all $x,y\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(x,\delta/2)\}\bigr|\;<\;C\,\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|,$
for any $C>1$. Hence $\mathcal{X}_{\delta}$ is a $(\delta,C)$–discretization of $K$. ∎
Next, we state a Lemma which we will use in the proof of Theorem 4.2.
Lemma 6.1.
Let $K$ be a compact metric space and $V\subset C(K)$ be compact. Given $\varepsilon>0$, there exists $\mathbf{y}:=\{y_{1},\cdots,y_{n}\}\in K$, and $\varphi_{1},\cdots,\varphi_{n}\in C(K)$ with $\|\varphi_{i}\|\leq 1$, such that
$\|f-\sum_{i=1}^{n}f(y_{i})\varphi_{i}\|_{C(K)}<\varepsilon$
for all $f\in V$. In particular, for any $f,g\in V$ we have
$\|f-g\|_{C(K)}\leq\|f-g\|_{\ell^{\infty}(\mathbf{y})}+2\varepsilon.$ (4)
Proof.
(of Lemma 6.1) Since $V$ is compact, by the Arzela-Ascoli theorem it is also equicontinuous, and so there exists $\delta>0$ such that if $|x-y|<\delta$, then $|f(x)-f(y)|<\varepsilon$ for all $f\in V$. Since $K$ is also compact, we can pick $y_{1},\cdots,y_{n}\in K$ such that $\{B(y_{i},\delta)\}_{i=1}^{n}$ form a covering of $K$. Finally, let $\{\varphi_{i}\}_{i=1}^{n}$ be a partition of unity subordinate to the balls $\{B(y_{i},\delta)\}_{i=1}^{n}$.
Given $y\in K$, for any $f\in V$, we have that,
$|f(y)-\sum_{i=1}^{n}f(y_{i})\varphi_{i}(y)|$ $=|\sum_{i=1}^{n}(f(y)-f(y_{i}))\varphi_{i}(y)|$
$<|\sum_{i,|y-y_{i}|<\delta}(f(y)-f(y_{i}))\cdot\varphi_{i}(y)|$
$<|\sum_{i,|y-y_{i}|<\delta}\varepsilon\cdot\varphi_{i}(y))|\leq\varepsilon.$
Using this, if $f,g\in V$ we have
$\begin{split}\|f-g\|_{C(K)}&\leq\|f-\sum_{i=1}^{n}f(y_{i})\varphi_{i}\|_{C(K)}+\|f-g\|_{\ell^{\infty}(\mathbf{y})}\max_{i}\|\varphi_{i}\|_{C(K)}+\|g-\sum_{i=1}^{n}g(y_{i})\varphi_{i}\|_{C(K)}\\ &\leq\|f-g\|_{\ell^{\infty}(\mathbf{y})}+2\varepsilon.\end{split}$ (5)
We can now return to the proof of Thm 4.2.
Proof.
(of Theorem 4.2) Fix $\varepsilon>0$. Since $\mathcal{K}$ is compact and $V$ is compact, it follows that the set
$\{G(u):~G\in\mathcal{K},~u\in V\}\subset C(K_{2})$ (6)
is compact. Since $K_{2}$ is also compact, we can apply Lemma 6.1 to get that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
$\|G(u_{q})-\sum_{i=1}^{n}G(u_{q})(y_{i})\gamma_{i}\|_{C(K_{2})}<\varepsilon/2,$
for every $G\in\mathcal{K}$ and $u_{q}\in V$. So let $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ be defined as,
$\Psi(y)=\{\gamma_{i}(y)\}_{i=1}^{n}.$
Using the notation $G(u_{q})(\textbf{y}):=(G(u_{q})(y_{1}),\cdots,G(u_{q})(y_{n}))$, we have that
$\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K_{2})}<\varepsilon/2.$ (7)
for every $G\in\mathcal{K}$.
We now show that $G(u_{q})(\textbf{y})$ can be approximated arbitrarily well by functions of the form
$\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right),$ (8)
where $\textbf{x}=(x_{1},...,x_{k})$ is a grid in $K_{1}$, $\Phi$ and $\rho$ are continuous, and $u_{1},...,u_{m}$ is an arbitrary $(\delta,C)$-discretization of $V$.
Since $\mathcal{K}$ is compact, it is equi-continuous, so there exists a $\delta$ such that if $\|u-v\|_{C(K_{1})}<25\delta/6$, then $\|G(u)-G(v)\|_{C(K_{2})}<\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$ for every $G\in\mathcal{K}$. By Lemma 6.1, there exists $\textbf{x}=\{x_{1},\cdots,x_{k}\}\in K_{1}$ such that
$\|u-v\|_{C(K_{1})}\leq\|u-v\|_{\ell^{\infty}(\mathbf{x})}+\frac{\delta}{6}.$ (9)
We let x be our grid in $K_{1}$ and $\delta$ be the value given in the Theorem.
Let $u_{1},...,u_{m}\in V$ be a $(\delta,C)$-discretization of $V$, and consider the restrictions $u_{i}(\mathbf{x})$ of $u_{i}$ to the grid $\mathbf{x}$, which lie in the compact set $V(\mathbf{x})\subset\mathbb{R}^{k}$. Let $u\in K$ and observe that by (9) we have
$\|u-u_{i}\|_{C(K_{1})}-\frac{\delta}{6}\leq\|u-u_{i}\|_{\ell^{\infty}(\mathbf{x})}\leq\|u-u_{i}\|_{C(K_{1})}$ (10)
This implies that
$\{u_{i}(\mathbf{x})\in B(u(\mathbf{x}),\delta/3)\}\subset\{u_{i}\in B(u,\delta/2)\}$ (11)
and that
$\{u_{i}\in B(u,\delta)\}\subset\{u_{i}(\mathbf{x})\in B(u(\mathbf{x}),\delta)\}.$ (12)
Hence, since $u_{1},..,u_{m}$ is a $(\delta,C)$-discretization, when restricted to the grid $\mathbf{x}$ we still have
$|\{u_{i}(\mathbf{x})\in B(u(\mathbf{x}),\delta/3)\}|\leq C|\{u_{i}(\mathbf{x})\in B(v(\mathbf{x}),\delta)\}|$ (13)
for any $u,v\in V$.
Next, we choose a maximal $2\delta$-separated subset $\widetilde{u}_{1}(\mathbf{x}),...,\widetilde{u}_{l}(\mathbf{x})$ of the set $V(\mathbf{x})\subset\mathbb{R}^{k}$, and let $\varphi_{j}:\mathbb{R}^{k}\rightarrow\mathbb{R}_{\geq 0}$ be a partition of unity subordinate to the covering $B(\widetilde{u}_{j}(\mathbf{x}),2\delta)$ of $V(\mathbf{x})$. Observe that the balls of radius $\delta$ about $\widetilde{u}_{j}(\mathbf{x})$ are disjoint, and so we have
$\varphi_{j}(v(\mathbf{x}))=1~~\text{whenever $\|v-\widetilde{u}_{j}\|_{\ell^{\infty}(\mathbf{x})}\leq\delta$.}$ (14)
In addition, by construction we have
$\varphi_{j}(v(\mathbf{x}))=0~~\text{whenever $\|v-\widetilde{u}_{j}\|_{\ell^{\infty}(\mathbf{x})}\geq 2\delta$.}$ (15)
Let $\Phi:\mathbb{R}^{k+n}\rightarrow\mathbb{R}^{l(n+1)}$ (so that $N=l(n+1)$) be defined by
$\Phi(\mathbf{u},\mathbf{v})=\{\varphi_{j}(\mathbf{u})\cdot\mathbf{v},\varphi_{j}(\mathbf{u})\}_{j=1}^{l}.$
We then have that for any $G\in\mathcal{K}$,
$\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))=\left\{\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\cdot G(u_{i})(\textbf{y}),\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\right\}_{j=1}^{l}$
Define $\xi:\mathbb{R}^{l(n+1)}\rightarrow\mathbb{R}^{ln}$ as,
$\xi(z_{1},\cdots,z_{l(n+1)})=\left\{\frac{(z_{j},z_{j+1},\cdots,z_{j+n-1})}{z_{j+n}}\right\}_{j\in[1,(n+1)+1,2(n+1)+1,\cdots,(l-1)(n+1)+1]}.$
Then, given the input from $\Phi$, we will have,
$\xi\left(\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)=\left\{\frac{\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\cdot G(u_{i})(\textbf{y})}{\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\right\}_{j=1}^{l}:=\{\overline{G(\widetilde{u}_{j})(\textbf{y})}\}_{j=1}^{l}.$
Here the $\overline{G(\widetilde{u}_{j})(\textbf{y})}$ notation is used just to emphasize that this is just the weighted average of the $G(u_{i})(\textbf{y})$ samples that are in the ball of radius $\delta$ centered at $\widetilde{u}_{j}$. We then define, $\rho:\mathbb{R}^{k+l(n+1)}\rightarrow\mathbb{R}^{n}$, as
$\rho(x,y)=\sum_{j=1}^{l}\textbf{(}\xi(y)\textbf{)}_{j}\cdot\varphi_{j}(x),$ (16)
so that we have
$\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)=\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x})).$ (17)
Now, fix any $G\in\mathcal{K}$. We will show that the above expression approximates $G(u_{q})(\textbf{y})$. We calculate
$\left\|G(u_{q})(\textbf{y})-\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}$ $=\left\|\sum_{j=1}^{l}G(u_{q})(\textbf{y})\cdot\varphi_{j}(u_{q}(\textbf{x}))-\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}$
$=\left\|\sum_{j,\|u_{q}-\widetilde{u_{j}}\|_{\ell^{\infty}(\textbf{x})}<2\delta}(G(u_{q})(\textbf{y})-\overline{G(\widetilde{u}_{j})(\textbf{y}))}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}$
Then, note
$\left\|G(u_{q})(\textbf{y})-\overline{G(\widetilde{u}_{j})(\textbf{y})}\right\|_{\ell^{\infty}(\textbf{y})}$ $=\left\|G(u_{q})(\textbf{y})-\frac{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\cdot G(u_{i})(\textbf{y})}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\right\|_{\ell^{\infty}(\textbf{y})}$
$=\left\|G(u_{q})(\textbf{y})-\sum_{i=1}^{m}\frac{\varphi_{j}(u_{i}(\textbf{x}))}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\cdot G(u_{i})(\textbf{y})\right\|_{\ell^{\infty}(\textbf{y})}$
$=\left\|\sum_{i=1}^{m}\frac{\varphi_{j}(u_{i}(\textbf{x}))}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\cdot\left(G(u_{q})(\textbf{y})-G(u_{i})(\textbf{y})\right)\right\|_{\ell^{\infty}(\textbf{y})}$
$\leq\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})\left|\sum_{i=1}^{m}\frac{\varphi_{j}(u_{i}(\textbf{x}))}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\right|=\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$
Where we use that $\|u_{q}-\widetilde{u}_{j}\|_{\ell^{\infty}(\textbf{x})}<2\delta$, $\|\widetilde{u}_{j}-u_{i}\|_{\ell^{\infty}(\textbf{x})}<2\delta$, and (9) imply that
$\|u_{q}-u_{i}\|_{C(K_{1})}\leq\|u_{q}-\widetilde{u}_{j}\|_{\ell^{\infty}(\textbf{x})}+\|\widetilde{u}_{j}-u_{i}\|_{\ell^{\infty}(\textbf{x})}+\delta/6<25\delta/6,$ (18)
so that by the choice of $\delta$, $\|G(u_{q})-G(u_{i})\|_{C(K_{2})}<\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$.
So we have
$\left\|G(u_{q})(\textbf{y})-\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}\leq\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})}),$
which combining with (7) and using (17) gives us,
$\|G(u_{q})-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
$\leq\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K)}+\|G(u_{q})(\textbf{y})\cdot\Psi-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
$<\varepsilon/2+\varepsilon/2=\varepsilon.$
Finally, we need to show that $\rho$ and $\Phi$ can be chosen continuous. It is clear from the definition that $\Phi$ is continuous. To show that $\rho$ is also continuous, we need to show that the denominator
$\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))$ (19)
in the definition of $\xi$ is bounded away from zero uniformly for any $(\delta,C)$-discretization $u_{1},...,u_{m}$ of $V$.
To prove this, we use (13) and (14). Let $B(v_{1}(\textbf{x}),\delta/3),...,B(v_{t}(\textbf{x}),\delta/3)$ be a covering of $V(\textbf{x})$. It follows that one of these balls must contain at least $m/t$ elements $u_{i}(\textbf{x})$ (recall that $m$ is the size of the $(\delta,C)$-discretization, which is not controlled). Using (13) this means each ball $B(u_{i}(\textbf{x}),\delta)$ must contain at least $m/(Ct)$ elements of the $(\delta,C)$-discretization. Finally, using (14) this gives
$\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\geq\frac{1}{m}\frac{m}{Ct}=\frac{1}{Ct}>0.$ (20)
This latter quantity is bounded independently of the $(\delta,C)$-discretization. This completes the proof. ∎

### v2

@@
 Abstract
-In-context Learning (ICL) is the remarkable capability displayed by some machine learning models to learn from examples in a prompt, without any further weight updates.
-ICL had originally been thought to emerge from the self-attention mechanism in autoregressive transformer architectures.
-DeepOSets is a non-autoregressive, non-attention based neural architecture that combines set learning via the DeepSets architecture with operator learning via Deep Operator Networks (DeepONets).
-In a previous study, DeepOSets was shown to display ICL capabilities in supervised learning problems.
-In this paper, we show that the DeepOSets architecture, with the appropriate modifications, is a multi-operator in-context learner that can recover the solution operator of a new PDE, not seen during training, from example pairs of parameter and solution placed in a user prompt, without any weight updates.
-Furthermore, we show that DeepOSets is a universal uniform approximator over a class of continuous operators, which we believe is the first result of its kind in the literature of scientific machine learning.
-This means that a single DeepOSets architecture exists that approximates in-context any continuous operator in the class to any fixed desired degree accuracy, given an appropriate number of examples in the prompt.
-Experiments with Poisson and reaction-diffusion forward and inverse boundary-value problems demonstrate the ability of the proposed model to use in-context examples to predict accurately the solutions corresponding to parameter queries for PDEs not seen during training.
+An important application of neural networks to scientific computing has been the learning of non-linear operators.
+In this framework, a neural network is trained to fit a non-linear map between two infinite dimensional spaces, for example, the solution operator of ordinary and partial differential equations.
+Recently, inspired by the discovery of in-context learning for large language models, an even more ambitious paradigm has been explored, called multi-operator learning.
+In this approach, a neural network is trained to learn many different operators at the same time.
+In order to evaluate one of the learned operators, the network is passed example inputs and outputs to disambiguate the desired operator.
+In this work, we provide a precise mathematical formulation of the multi-operator learning problem.
+In addition, we modify a simple efficient architecture, called DeepOSets, for multi-operator learning and prove its universality for multi-operator learning.
+Finally, we provide a comprehensive set of experiments that demonstrate the ability of DeepOSets to learn multiple operators corresponding to different initial-value and boundary-value differential equations and use in-context examples to predict accurately the solutions corresponding to queries and differential equations not seen during training.
+The main advantage of DeepOSets is its architectural simplicity, which allows the derivation of theoretical guarantees and training times that are in the order of minutes, in contrast to similar transformer-based alternatives that are empirically justified and require hours of training.
+**footnotetext: These authors contributed equally to this work.
 1 Introduction
-In learning the solution operator of a parametric PDE from sample data, one is not interested in merely obtaining a single solution for a single PDE, but to learn the mapping between parameter and solution, so that new solutions corresponding to new values of the parameter can be obtained quickly and without further training.
+Learning non-linear operators from data using neural networks has recently proved to be a foundational application of deep learning to scientific computing ((Lu et al., 2021; Li et al., 2020; Boullé and Townsend, 2024).
+In learning the solution operator of a parametric PDE from sample data, one is not interested in merely obtaining a single solution for a single PDE (as is typical in the popular Physics-informed neural network approach ((Raissi et al., 2019)), but instead learning the mapping between parameter and solution, so that new solutions corresponding to new parameter values can be obtained quickly and without further training.
 Furthermore, both the parameter and the solution are typically functions in infinite-dimensional spaces.
 For example, the problem might be to learn the space-varying fluid pressure (solution) corresponding to the space-varying permeability in a porous medium (parameter), where the physical process is governed by Darcy’s Equation.
-The training data in this case consists of pairs of parameter and corresponding solution, obtained from experiments or, more commonly, by solving Darcy’s equation using a high-fidelity solver.
-The trained operator allows quick exploration of the parameter space, which is unfeasible to do with expensive experiments or high-fidelity solves.
-The most well-known operator learners are DeepONets ludeeponet2021 and Fourier Neural Operators li2020fourier, which are both based on specific deep neural network architectures.
-A more recent development is foundation models for scientific problems, which is inspired by the success of of large language models.
-A foundation model can be trained on a large body of data and be adapted for new tasks, perhaps with no further training bommasani2021opportunities.
-In the context of PDEs, foundation models correspond to multiple operator (multi-operator) learning.
-The challenge could be to be able to train a model with data from diverse PDEs with diverse equation coefficients and boundary conditions, and then produce fast solutions for new PDEs on demand.
+The training data in this case consists of pairs of parameters and corresponding solutions, obtained from experiments or, more commonly, by solving Darcy’s equation using a high-fidelity solver.
+The trained operator allows quick exploration of the parameter space, which is infeasible to do with expensive experiments or high-fidelity solvers.
+The most well-known operator learners are DeepONets ((Lu et al., 2021) and Fourier Neural Operators ((Li et al., 2020), which are both based on specific deep neural network architectures.
+As in the case of standard multi-layer perceptrons, the theoretical foundations of this approach are a universal approximation theorem, originally due to Chen and Chen ((1995b), which show that a precursor to the DeepONet ((Lu et al., 2021) architecture is universal.
+Specifically, this means that DeepONet can in principle approximate to arbitrary accuracy any continuous non-linear map $F:K\rightarrow C(K_{2})$, where $K_{1}\subset\mathbb{R}^{d}$ and $K_{2}\subset\mathbb{R}^{n}$ are compact subsets and $K\subset C(K_{1})$ is a compact subset of $C(K_{1})$.
+A more recent development is foundation models for scientific problems, which is inspired by the success of large language models.
+A foundation model can be trained on a large body of data and be adapted for new tasks, perhaps with no further training ((Bommasani, 2021).
+In the context of differential equations, foundation models correspond to multiple operator (multi-operator) learning.
+The challenge here could be to be able to train a model with data from diverse differential equations with diverse equation coefficients and boundary conditions, and then produce fast solutions for new PDEs on demand.
 In contrast to single operator learning, here a well-posedness question arises, since multiple solution operators acting on the same parameter function will produce different solutions.
 How is the model supposed to identify the correct operator?
-There are two basic methods to resolve this issue: the first, is to fine-tune the model on a small amount of data from the new PDE; the second is to use in-context learning (ICL), when the PDE is identified from example pairs in a prompt with no further training.
-In-Context Operator Networks (ICON) yang2023context and its subsequent variants yang2024pde; yang2023ICON-LM; zhang2025GenICON remain the only published approach to in-context learning of PDE solution operators (see Section 2 for other work related to multi-operator learning and foundation models for PDEs).
-ICON is a framework that adapts GPT-style in-context learning to PDE problems, by employing a decoder transformer to generate the solution from a sequential prompt followed by the query.
-ICON is thus an autoregressive, attention-based approach, with quadratic complexity in prompt length, due to the dot-product self-attention mechanism.
+There are two basic methods to resolve this issue: the first, is to fine-tune the model on a small amount of data from the new equation; the second is to use in-context learning (ICL), when the equation is identified from example pairs in a prompt with no further training.
+In-Context Operator Networks (ICON) ((Yang et al., 2023a) and its subsequent variants ((Yang and Osher, 2024; Yang et al., 2023c; Zhang et al., 2025) is the best-known approach to in-context learning of PDE solution operators (see Section 2 for other work related to multi-operator learning and foundation models for PDEs).
+ICON is a framework that adapts GPT-style in-context learning to differential equation problems, by employing a decoder transformer to generate the solution from a sequential prompt followed by the query.
+ICON is thus an attention-based approach, with quadratic complexity in prompt length, due to the dot-product self-attention mechanism.
 This can introduce significant computational overhead and loss of accuracy with very long prompts.
-In addition, if the query changes, the prompt has to be processed sequentially again from the beginning (unless caching is used).
-In this paper, we propose an alternative approach to in-context learning of PDE solution operators, which processes the prompt non-autoregressively, in a fully parallel manner, without using self-attention, being therefore much faster during both training and inference time.
-This approach is based on DeepOSets, a neural architecture that combines set learning via the DeepSets architecture zaheer2017deep with operator learning via the previously mentioned DeepONets architecture.
+In addition, if the query changes, the prompt has to be processed again (unless caching is used).
+In addition, a theoretical study of in-context multi-operator learning remains incomplete.
+In particular, it is not known whether existing architectures such as ICON are universal approximators.
+One key difficulty is that the lack of a precise mathematical formulation of the multi-operator learning problem, i.e., what does it mean to simultaneously approximate multiple operators at the same time?
+In this paper, we provide a precise mathematical formulation of the in-context multi-operator learning problem, which is formulated as learning a compact set of operators between two Banach spaces, instead of just a single operator.
+Further, we propose an alternative approach to in-context learning of PDE solution operators, which processes the prompt non-autoregressively, in a fully parallel manner, without using self-attention, being therefore much faster during both training and inference time.
+This approach is based on DeepOSets, a neural architecture that combines set learning via the DeepSets architecture ((Zaheer et al., 2017) with operator learning via the previously mentioned DeepONets architecture.
 In DeepOSets, the DeepSets layer acts as an encoder of the prompt into a latent representation that is the input, along with the query, to the branch path of a DeepONet.
 The DeepSets layer confers linear complexity in the length of the prompt, in contrast to the quadratic complexity of self-attention, and a built-in permutation-invariance inductive bias to the order of the examples in the prompt.
-DeepOSets was shown in a previous publication to display ICL capabilities in regression problems chiu2024deeposets.
-Here, we introduce architectural modifications to DeepOSets that allow it to act as a multi-operator learner for parametric PDEs.
+DeepOSets has been shown to display ICL capabilities in regression problems ((Chiu et al., 2024).
+Here, we introduce architectural modifications to DeepOSets that allow it to act as a multi-operator learner for parametric differential equations.
 This is accomplished by discretizing a variable number of solution and parameter function pairs, which are entered into the prompt and encoded by the DeepSets layer into latent representation.
 The latter is concatenated to the discretization of the query parameter function and entered directly into the branch path of the DeepONet.
-Crucially, we prove that this DeepOSets architecture approximates uniformly a compact set of continuous solution operators, which means that DeepOSets are not only universal approximators of single solution operators, like DeepONets, but they are universal approximators of multiple operators, in the sense that a single DeepOSets architecture exists that can approximate any operator in the class to the same fixed degree of accuracy.
-This guarantees that DeepOSets can learn in-context the solution operators of multiple PDEs.
-Experiments with Poisson and reaction-diffusion forward and inverse problems demonstrate that DeepOSets can accurately predict the solution to a new PDE with inverse or forward problem and unknown value of equation coefficients and boundary conditions learned in-context.
+A detailed description of our architecture can be found in Section 4 and Figure 1.
+Crucially, we prove that this DeepOSets architecture is universal for the in-context multi-operator learning problem.
+Specifically, we prove that it can approximate uniformly any compact set of continuous solution operators.
+This means that DeepOSets are not only universal approximators of single solution operators, like DeepONet, but they are universal approximators of multiple operators, in the sense that a single DeepOSets architecture exists that can approximate any operator in the class to the same fixed degree of accuracy.
+This guarantees that DeepOSets can learn in-context the solution operators of multiple differential equations.
+Finally, experiments with Poisson and reaction-diffusion forward and inverse problems demonstrate that DeepOSets can accurately predict the solution to a new differential equation with inverse or forward versions and unknown value of equation coefficients and boundary conditions learned in-context.
+Training times are in the order of minutes, as compared to hours for similar transformer-based alternatives.
+1.1 Our Contributions
+To summarize, our main contributions are:
+  1. 1.
+    We provide a precise mathematical formulation of the in-context multi-operator learning problem and universal approximation problem in Section 3 and Definition 3.3, respectively.
+  2. 2.
+    We prove that a variant of DeepOSets, defined in Section 4, is universal for in-context multi-operator learning within Section 5 in Theorem 5.1.
+This is to the best of our knowledge the first universal approximation result for in-context multi-operator learning.
+  3. 3.
+    We demonstrate in Section 6 that DeepOSets is capable of jointly learning multiple operators and exhibits robust generalization to support sets larger than those encountered during training.
+Furthermore, training is very efficient, taking only minutes on a consumer GPU.
 2 Related Work
-MIONet is probably the first approach to attempt multi-input operator learning, by encoding several input functions via multiple branch network (and a shared trunk network) of a DeepONet, though it still learns a single operator jin2022mionet.
-POSEIDON is a large foundation model for PDEs based on a scalable Operator Transformer (scOT) with a vision transformer architecture, is pretrained on compressible Euler and incompressible Navier-Stokes equations and applied to 15 downstream tasks, showing better sample efficiency and accuracy than baselines herde2024poseidon.
-PROSE-PDE is a transformer-based multi-operator learning approach trained on one-dimensional time-dependent nonlinear PDEs that encodes both numerical and symbolic data types for forward and inverse PDE problems sun2025towards.
-LeMON is based on pretraining and fine-tuning for multi-operator learning, and introduces low-rank adaptation to reduce computational cost sun2024lemon.
-MT-DeepONet introduces modifications to the branch network of a DeepONet to accommodate various functional forms of parameterized coefficients and multiple geometries kumar2025synergistic.
-In-context learning for PDE solution operators was introduced by ICON yang2023context.
-The follow-up study on 1D scalar conservation laws provides a detailed methodology for applying ICON to temporally evolving PDEs yang2024pde, while ICON-LM integrates ICON concepts with language-model architectures yang2023ICON-LM.
-GenICON introduces a generative, probabilistic approach to ICON zhang2025GenICON.
+The universal approximation theorem for multi-layer perceptrons, which forms the basis for more sophisticated universal approximation theorems, was proved in Cybenko ((1989) and Hornik ((1991).
+Universal approximation of operators using neural networks was first proved in Chen and Chen ((1995b), following previous results from Chen and Chen ((1993, 1995a).
+This result was used to design the DeepONet operator learning architecture ((Lu et al., 2021).
+Subsequently, various universal architectures for operator learning have been proposed, a notable example being Fourier Neural Operators ((Li et al., 2020).
+These architectures typically evaluate the input function on a fixed grid, and an architecture compatible with variable input grids was proposed and analyzed in Prasthofer et al.
+((2022).
+In-context learning (ICL) is a phenomenon first observed in large language models ((Brown et al., 2020; Tsimpoukelli et al., 2021; Liu et al., 2023; Garg et al., 2022), whereby these models are able to learn new tasks without fine-tuning from examples presented to them in the prompt.
+This has sparked substantial theoretical research effort into the mechanism underlying this phenomenon.
+Transformers have been theoretically shown to exhibit ICL on a variety of linear models, including least squares, ridge regression, Lasso, generalized linear models and linear inverse problems ((Bai et al., 2023; Zhang et al., 2024; Von Oswald et al., 2023).
+Generalization error for transformer-based ICL for data lying on a low-dimensional manifold has been studied in ((Shen et al., 2025).
+In terms of applying ICL to operator learning, MIONet is to the best of our knowledge the first approach to attempt multi-input operator learning, by encoding several input functions via multiple branch network (and a shared trunk network) of a DeepONet, though it still learns a single operator ((Jin et al., 2022).
+POSEIDON is a large foundation model for PDEs based on a scalable Operator Transformer (scOT) with a vision transformer architecture, is pretrained on compressible Euler and incompressible Navier-Stokes equations and applied to 15 downstream tasks, showing better sample efficiency and accuracy than baselines ((Herde et al., 2024).
+PROSE is a transformer-based multi-operator learning approach trained on one-dimensional time-dependent nonlinear PDEs that encodes both numerical and symbolic data types for forward and inverse PDE problems ((Sun et al., 2025).
+PROSE is used to build a foundation model for fluid dynamics problems in Liu et al.
+((2024).
+Recently, the PROSE approach was extended to include physics-informed losses ((Zhu et al., 2025).
+LeMON is based on pretraining and fine-tuning for multi-operator learning, and introduces low-rank adaptation to reduce computational cost ((Sun et al., 2024).
+MT-DeepONet introduces modifications to the branch network of a DeepONet to accommodate various functional forms of parameterized coefficients and multiple geometries ((Kumar et al., 2025).
+In-context learning for differential equation solution operators was introduced by ICON ((Yang et al., 2023a), with a follow-up study on 1D scalar conservation laws providing a detailed methodology for applying ICON to time-evolution PDEs ((Yang and Osher, 2024).
+ICON-LM integrates ICON concepts with language-model architectures ((Yang et al., 2023c).
+GenICON introduces a generative, probabilistic approach to ICON ((Zhang et al., 2025).
+In-context learning of linear PDEs using transformers has also been studied in ((Cole et al., 2024).
+3 Mathematical Formulation of In-context Operator Learning
+In this section, we give a precise mathematical formulation of the in-context multi-operator learning problem.
+Based upon this, we will be able to argue whether a given architecture is universal for this problem.
+Let $K_{1},K_{2}\subset\mathbb{R}^{d}$ be compact subsets of some Euclidean space.
+We will consider continuous maps between the Banach spaces $X:=C(K_{1})$ and $Y:=C(K_{2})$, whose domain is a compact subset $V\subset X$.
+Note that the dimensionality of $K_{1}$ and $K_{2}$ need not be the same in general, but to simplify the presentation we restrict to this case in the following.
+The minor adjustments required to handle the more general case where $K_{1}\subset\mathbb{R}^{d_{1}}$ and $K_{2}\subset\mathbb{R}^{d_{2}}$ are left to the reader.
+In traditional operator learning, a continuous function $G:V\rightarrow Y$ is approximated using a neural network.
+The input function $u\in V$ is discretized at a fixed set of points $x_{1},...,x_{k}\in K_{1}$, and the values $u(x_{1}),...,u(x_{k})$ together with an evaluation point $y\in K_{2}$ are passed to a neural network whose output approximates $G(u)(y)$.
+Universality in this context means that for any $G$ and $\epsilon>0$, there exists a grid $\textbf{x}:=\{x_{1},...,x_{k}\}\in K_{1}$ (depending on $G$ and $\epsilon$) and a neural network $G_{\epsilon}$ such that
+$\left|G_{\epsilon}(u(\textbf{x}),y)-G(u)(y)\right|<\varepsilon$
+for any $u\in V$ and $y\in K_{2}$ ((Chen and Chen, 1995b; Lu et al., 2021). Here we denote by $u(\textbf{x}):=\left(u(x_{i})\right)_{i=1}^{k}\in\mathbb{R}^{k}$ the function evaluated at the grid x.
+Instead of learning a single operator $G:V\rightarrow Y$, in-context multi-operator learning considers an entire class of operators. To formulate this, we consider the space of continuous operators
+$\mathcal{O}:=\{G:V\rightarrow Y~\text{continuous}\}$
+equipped with the induced operator norm given by
+$\|G\|_{\mathcal{O}}:=\sup_{x\in V}{\|G(x)\|_{Y}}.$
+We let $\mathcal{K}\subset\mathcal{O}$ be a compact subset of this space of operators.
+$\mathcal{K}$ is the class of operators that we wish to learn.
+In order to disambiguate the desired operator $G\in\mathcal{K}$, we pass the network a sequence of in-context examples $(u_{1},G(u_{1})),...,(u_{m},G(u_{m}))$ in addition to the input function $u\in K$ and the evaluation point $y\in K_{2}$.
+All of the functions $_{i}, uu$, and $G(u_{i})$ are discretized on fixed grids $\textbf{x}=(x_{1},...,x_{k})$ and $\textbf{y}=(y_{1},...,y_{n})$ in $K_{1}$ and $K_{2}$, respectively.
+We remark that one could additionally consider variable grids in $K_{1}$ and $K_{2}$, but we leave this to future work.
+Although the grids x and y are fixed, we want to allow the in-context examples to be given for arbitrary functions $u_{1},...,u_{m}\in V$ (even the number of functions $m$ should also be arbitrary).
+Of course, we need to make some assumption which guarantees that the $u_{i}$ sufficiently cover the domain $V$ so that the desired operator $G\in\mathcal{K}$ can be determined to the desired accuracy from the outputs $G(u_{i})$.
+For technical reasons, we also need to choose the $u_{i}$ to be sufficiently uniformly distributed within $V$.
+This is captured in the following definition.
+Definition 3.1 ($(\delta,C)$-discretization).
+Let $\delta>0$ and $C>1$. A $(\delta,C)$-discretization of a compact metric space $V$ is a set of points $\textbf{u}=\{u_{1},\cdots,u_{m}\}$ such that
+$|\{u_{i}\in B(u,\delta/2)\}|<C|\{u_{i}\in B(v,\delta)\}|\,$,
+for any $u,v\in V$.
+Intuitively, a $(\delta,C)$-discretization is a $\delta$-net of $V$ which is not too unbalanced, in the sense that the density of points is relatively uniform.
+Let us first note that this definition is not vacuous.
+Lemma 3.2.
+Let $V$ be a compact set in a metric space, $\delta>0$ and $C>1$. Then there always exists a $(\delta,C)$-discretization of $V$.
+Proof.
+See the Appendix.∎
+We can now formulate the concept of universality for in-context multi-operator learning.
+Definition 3.3 (Mathematical formulation of In-Context Operator Learning).
+We define universality for in-context operator learning to mean that for any $\mathcal{K}\subset\mathcal{O}$, any $\epsilon>0$ and $C>1$, there exist grids $\textbf{x}=(x_{1},...,x_{k})$ and $\textbf{y}=(y_{1},...,y_{n})$ in $K_{1}$ and $K_{2}$, respectively, a $\delta>0$ and a network $K_{\epsilon}$ such that
+$|K_{\epsilon}(u_{1}(\textbf{x}),G(u_{1})(\textbf{y}),...,u_{m}(\textbf{x}),G(u_{m})(\textbf{y}),u(\textbf{x}),y)-G(u)(y)|<\epsilon\,$, (1)
+for any $G\in\mathcal{K}$, $u\in V$, $y\in K_{2}$, and $(\delta,C)$-discretization $\textbf{u}=(u_{1},...,u_{m})$ of $V$.
+In order words, any operator of $\mathcal{K}$ can be approximated to arbitrary accuracy provided that the in-context examples are sufficiently dense in $V$.
+Note that the network architecture should be able to process an arbitrary, i.e., not a priori specified, number of in-context examples.
+This can be achieved using a transformer architecture ((Vaswani et al., 2017), for example.
+In the next section, we describe a simple architecture based upon Deep Sets ((Zaheer et al., 2017), called DeepOSets ((Chiu et al., 2024) for this purpose.
 4 The DeepOSets Architecture for In-Context Operator Learning
-The DeepOSets model is a hybrid neural network comprising a DeepSets encoder and a DeepONet decoder, as seen in Figure 1.
-The DeepSets module encodes non-autoregressively the discretized parameter and solution function example pairs in the prompt.
+The DeepOSets model is a hybrid neural network comprising of a DeepSets encoder and a DeepONet decoder, as seen in Figure 1.
+Figure 1: DeepOSets architecture for in-context multi-operator learning.
+The DeepSets module encodes the discretized parameter and solution function example pairs in the prompt.
 Each pair $(u_{i},G(u_{i}))$ is first mapped to an individual fixed-length embedding $\bf{h}_{i}$ by a shared multi-layer perceptron (MLP) denoted $\Phi$.
@@
 The output is therefore entirely mesh-free.
-Figure 1: DeepOSets architecture for in-context multi-operator learning.
+We denote the branch network as $\rho$, and the trunk network as $\Psi$.
+With this notation, the output of the DeepOSets architecture using mean pooling, given input $x_{q}$, can be mathematically expressed as
+$\hat{G}(u_{q})(x_{q}):=\Psi(x_{q})\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right).$ (2)
 5 Universality of DeepOSets for In-Context Operator Learning
-In this section, we prove that DeepOSets are universal uniform approximators over a compact set of continuous operators, i.e., a single DeepOSets neural network exists that bounds the supremum of the error over the entire class to any desired fixed accuracy.
+In this section, we prove that DeepOSets is a universal uniform approximator for any compact set of continuous operators in the sense of Definition 3.3, i.e., a single DeepOSets neural network exists that bounds the supremum of the error over the entire class to any desired fixed accuracy.
 This shows that a unique DeepOSets neural network could be learned that approximates in-context any continuous operator to the desired accuracy, given an appropriate number of examples in the prompt.
-We first define the class of operators of interest.
-Let $X:=C(K_{1}),~Y:=C(K_{2})$, with $K_{1},K_{2}\subset\mathbb{R}^{d}$ compact subsets of some Euclidean space.
-Let $V\subset X$ be compact.
-Given this, consider the space of continuous operators (in the induced operator norm):
-$\mathcal{O}:=\{G:V\rightarrow Y~\text{continuous}\}$
-with norm given by
-$\|G\|_{\mathcal{O}}:=\sup_{x\in V}{\|G(x)\|_{Y}}.$
-We let $\mathcal{K}\subset\mathcal{O}$ to be compact subset of this space of operators.
-We will need to discretize functions in $X$ on different grids in $K_{1}$ and will also need to consider certain coverings of the set $V$.
-For this purpose, we introduce the following definition.
-Definition 4.1 (Grid).
-Let $\delta>0$ and $C>1$. A $(\delta,C)$-discretization of a compact metric space $K$ is a set of points $\textbf{x}=\{x_{1},\cdots,x_{n}\}$ such that
-$|\{x_{i}\in B(x,\delta/2)\}|<C|\{x_{i}\in B(y,\delta)\}|$
-for any $x,y\in K$.
-Intuitively, a $(\delta,C)$-discretization is a $\delta$-net of $K$ which is not too unbalanced, in the sense that the density of points is relatively uniform.
-Let us first note that this definition is not vacuous.
-Lemma 4.1.
-Let $K$ be a compact set in a metric space, $\delta>0$ and $C>1$. Then there always exists a $(\delta,C)$-discretization of $K$.
+Theorem 5.1 (Uniform universality for In-Context Operator Learning).
+Given $\varepsilon>0$, $C>1$ and a compact set $\mathcal{K}\subset\mathcal{O}$, there exist integers $N,M\geq 1$, $\delta>0$, fixed discretizations $\textbf{x}=\{x_{1},...,x_{k}\}$ of $K_{1}$ and $\textbf{y}=\{y_{1},...,y_{n}\}$ of $K_{2}$, and continuous functions $\Psi:\mathbb{R}^{d}\!\rightarrow\mathbb{R}^{M}$, $\rho:\mathbb{R}^{k+N}\!\rightarrow\mathbb{R}^{M}$, $\Phi:\mathbb{R}^{k+n}\!\rightarrow\mathbb{R}^{N}$, such that
+$\sup_{G\in\mathcal{K}}\>\Bigg|\,G(u_{q})(x_{q})~-~\left.\Psi(x_{q})\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\right|<\varepsilon\,$, (3)
+for any $(\delta,C)$-discretization $\textbf{u}:=(u_{1},\cdots,u_{m})$ of $V$, any $u_{q}\in V$, and any $x_{q}\in K_{2}$.
 Proof.
-See the Appendix.∎
-Given a continuous function $u:K\rightarrow\mathbb{R}$ and a $(\delta,C)$-discretization $\textbf{x}=(x_{1},...,x_{n})$, we denote by
-$u(\textbf{x}):=\left(u(x_{i})\right)_{i=1}^{n}\in\mathbb{R}^{n}$ (1)
-for the function evaluated at the grid x. Given these definitions and notation, our main result is the following.
-Theorem 4.2 (Uniform universality for In-Context Operator Learning).
-Given $\varepsilon>0$, $C>1$ and a compact set $\mathcal{K}\subset\mathcal{O}$, there exist integers $N,M\geq 1$, $\delta>0$, fixed discretizations $\textbf{x}=\{x_{1},...,x_{k}\}$ of $K_{1}$ and $\{y_{1},...,y_{n}\}$ of $K_{2}$, and continuous functions $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{M}$, $\rho:\mathbb{R}^{k+N}\rightarrow\mathbb{R}^{M}$, and $\Phi:\mathbb{R}^{k+n}\rightarrow\mathbb{R}^{N}$, such that
-$\sup_{G\in\mathcal{K}}\>\left|\,G(u_{q})(y)-\Psi(y)\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\right|<\varepsilon.$
-for any $(\delta,C)$-discretization $\textbf{u}:=(u_{1},\cdots,u_{m})$ of $V$, any $u_{q}\in V$, and any $y\in K_{2}$.
-Proof.
-See the Appendix.∎
-The continuous functions guaranteed by Theorem 4.2 can be approximated using neural networks by invoking standard universal approximation theorems for MLPs cybenko1989approximation; hornik1991approximation; pinkus1999approximation.
+For the full detailed proof, please see the Appendix.
+Here, we present an overview of the proof, in which the approximation error is decomposed into two inequalities.
+The compactness assumption enables this decomposition to hold uniformly over the class of continuous operators by reducing the problem to finite-dimensional approximations via discretizations on the grids $\mathbf{x}$ and $\mathbf{y}$.
+For the first inequality, using a partition of unity Lemma (in Appendix), we show that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
+$\|G(u_{q})-\sum_{i=1}^{n}G(u_{q})(y_{i})\gamma_{i}\|_{C(K_{2})}<\varepsilon/2,$
+for every $G\in\mathcal{K}$ and $u_{q}\in V$.
+For the second inequality, we show that there exists a grid x of $K_{1}$ such that
+$G(u_{q})(\textbf{y}):=(G(u_{q})(y_{1}),\cdots,G(u_{q})(y_{n}))$
+can be approximated well by functions of the form,
+$\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right),$ (4)
+where $u_{1},...,u_{m}$ is a $(\delta,C)$-discretization of $V$.
+We construct $\Phi$ using a partition-of-unity argument to be localized/subordinate over a discretization of the input space, and define $\rho$ by forming normalized local averages of the corresponding operator output samples evaluated on the output grid and recombining them into a continuous finite-dimensional approximation.
+We then let $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$, (here $n=M)$, be defined as,
+$\Psi(x)=\{\gamma_{i}(x)\}_{i=1}^{n}\,.$
+Taking the dot product with (4) and applying the triangle inequality together with the two bounds yields the desired approximation inequality.
+Finally, we verify that the constructions $\Phi$ and $\rho$ are continuous.
+Continuity of $\Phi$ is immediate from its definition, while continuity of $\rho$ follows from uniform lower bounds on the normalization factors ensured by the balanced $(\delta,C)$-discretization of $V$.
+∎
+The continuous functions guaranteed by Theorem 5.1 can be approximated using neural networks by invoking standard universal approximation theorems for MLPs ((Cybenko, 1989; Hornik, 1991; Pinkus, 1999).
 Here it is important that we are averaging the outputs of $\Phi$ to ensure that the function $\rho$ only needs to be approximated on a compact set.
 6 Experimental Results
-Here, we demonstrate that the proposed DeepOSets architecture can learn the solution operator for multiple PDEs, with the correct PDE being disambiguated with the prompt examples, without any further weight updates.
-We consider two one-dimensional parametric PDEs and forward and inverse problems for each.
-The Poisson boundary-value PDE reads:
-$\frac{d^{2}u(x)}{dx^{2}}$ $=\,f(x)\,,\quad 0<x<1\,$, (2)
-with $u(0)=u_{0}$ and $u(1)=u_{1}$.
-The forward problem is to find the solution $u(x)$ given the forcing function $f(x)$.
-The inverse or control problem is to find the forcing function $f(x)$ that produces the given trajectory $u(x)$.
-Very importantly, the boundary values $u_{0}$ and $u_{1}$ are unknown and must be learned in-context.
-The reaction-diffusion boundary-value PDE is:
-$a\frac{d^{2}u(x)}{dx^{2}}+k(x)u(x)=c\,,\quad 0<x<1\,$, (3)
-with $u(0)=u_{0}$ and $u(1)=u_{1}$.
-The forward problem is to find the solution $u(x)$ given the coefficient function $k(x)$.
-The inverse problem is to find the coefficient function $k(x)$ that produces the given solution $u(x)$.
-Here, the coefficients $a$ and $c$ and the boundary values $u_{0}$ and $u_{1}$ are all unknown and must be learned in-context.
-(a)         Forward Poisson
-    $\frac{d^{2}{\color[rgb]{1,0,0}G(u)(x)}}{dx^{2}}={\color[rgb]{0,0,1}u(x)}$, given $u(0),u(1)$
-(b)         Inverse Poisson
-    $\frac{d^{2}{\color[rgb]{0,0,1}u(x)}}{dx^{2}}={\color[rgb]{1,0,0}G(u)(x)}$
-, given $u(0),u(1)$
-(c)      Forward Reaction-Diffusion
-   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{1,0,0}u(x)}+{\color[rgb]{0,0,1}k(x)}{\color[rgb]{1,0,0}u(x)}=c$, given $a$, $c$
-(d)      Inverse Reaction-Diffusion
-   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{0,0,1}u(x)}+{\color[rgb]{1,0,0}k(x)}{\color[rgb]{0,0,1}u(x)}=c$, given $a$, $c$
-Figure 2: In-context PDE solution operator learning with the DeepOSets architecture.
-In each case, the top plot contains the parameter functions while the bottom plot contains the corresponding solution functions.
-The gray lines indicate in each case the parameter-solution function pairs given in the prompt.
-The blue line indicates the query parameter function, while the red line indicates the corresponding solution predicted by the trained DeepOSets model.
-The black line indicates the exact solution corresponding to the query, using the appropriate PDE.
-The DeepOsets model is trained to learn the four PDE problems simultaneously, with coefficients sampled randomly.
-All PDE coefficients and boundary values are chosen randomly, so that the number of solution operators involved is in fact infinite.
+# Problem Description Differential Equations Parameters Conditions QoIs
+1 Forward IVP 1 $\frac{d}{dt}u(t)=a_{1}c(t)+a_{2}$ for $t\in[0,1]$ $a_{1},a_{2}$ $u(0),c(t),t\in[0,1]$ $u(t),t\in[0,1]$
+2 Inverse IVP 1 $u(t),t\in[0,1]$ $c(t),t\in[0,1]$
+3 Forward IVP 2 $\frac{d}{dt}u(t)=a_{1}c(t)u(t)+a_{2}$ for $t\in[0,1]$ $a_{1},a_{2}$ $u(0),c(t),t\in[0,1]$ $u(t),t\in[0,1]$
+4 Inverse IVP 2 $u(t),t\in[0,1]$ $c(t),t\in[0,1]$
+5 Forward IVP 3 $\frac{d}{dt}u(t)=a_{1}u(t)+a_{2}c(t)+a_{3}$ for $t\in[0,1]$ $a_{1},a_{2},a_{3}$ $u(0),c(t),t\in[0,1]$ $u(t),t\in[0,1]$
+6 Inverse IVP 3 $u(t),t\in[0,1]$ $c(t),t\in[0,1]$
+7 Forward Poisson BVP $\frac{d^{2}}{dx^{2}}u(x)=f(x)$ for $x\in[0,1]$ $u(0),u(1)$ $f(x),x\in[0,1]$ $u(x),x\in[0,1]$
+8 Inverse Poisson BVP $u(x),x\in[0,1]$ $f(x),x\in[0,1]$
+9 Forward linear reaction-diffusion BVP $a\frac{d^{2}}{dx^{2}}u(x)+k(x)u(x)=c$ for $x\in[0,1]$ $u(0),u(1),a,c$ $k(x),x\in[0,1]$ $u(x),x\in[0,1]$
+10 Inverse linear reaction-diffusion BVP $u(x),x\in[0,1]$ $k(x),x\in[0,1]$
+11 Forward nonlinear reaction-diffusion BVP $a\frac{d^{2}}{dx^{2}}u(x)+k(x)u(x)^{3}=c$ for $x\in[0,1]$ $u(0),u(1),a,c$ $k(x),x\in[0,1]$ $u(x),x\in[0,1]$
+12 Inverse nonlinear reaction-diffusion BVP $u(x),x\in[0,1]$ $k(x),x\in[0,1]$
+Table 1: Summary of tasks.
+Two benchmarks are considered: in the first, only the first six tasks, corresponding to the initial-value problems (IVP), are considered; in the second, all twelve tasks, consisting of both IVPs and boundary-value problems (BVP), are considered jointly.
+We demonstrate that the DeepOSets architecture successfully learns the solution operator for various initial-value problems (IVP) and boundary-value problems (BVP).
+Our evaluation spans 12 distinct tasks, which are summarized in Table 1.
+This table and the terminology “condition” and “quantity of interest (QoI)” are adapted from Yang et al.
+((2023b).
+There are in fact an infinite number of solution operators in this experiment, corresponding to different settings of equation coefficients and boundary values.
+The prompt examples allow the model to disambiguate the specific operator in-context, without requiring additional weight updates.
+We fix $u(0)=0$ in all IVPs (tasks 1–6).
+In the BVPs (tasks 7–12), the boundary values $u(0)$ and $u(1)$ are chosen randomly and are unknown to the model.
+To evaluate how task range and diversity influence performance, we consider two benchmarks: in the first, the DeepOSets model is trained and tested only on the 6 boundary-value problems, while in the second, the model is trained and tested on all 12 initial and boundary value problems.
+As will be seen below, training on the 12 tasks indeed produces a more accurate multi-operator learner.
+In all experiments, equation coefficients and boundary condition values are chosen randomly.
 The physical domain in all cases is the interval $[0,1]$, which is discretized uniformly into 100 points.
-The prompts during both training and testing contain 4 example pairs of parameter and solution functions.
-The MLP $\phi$ in the DeepOSets contains 2 hidden layers of 400 neurons each, and the latent vector encoding the prompt data has size 500.
+The MLP $\phi$ in the DeepOSets contains 2 hidden layers of 500 neurons each, and the latent vector encoding the prompt data has size 500.
 The pooling operation is the sample mean.
-The branch and trunk networks in the DeepOSets contain hidden layers of 200, 400, and 200 neurons with intermediate layer of 150 neurons.
+The branch and trunk networks in the DeepOSets model contain hidden layers of 200, 500, and 200 neurons.
+The pooled latent space is 2000 with mean pooling.
 All neural network use ReLU as the activation function.
-To generate the training data, for each of the four problems (Poisson/reaction-fiffusion, forward/inverse), 100 parameter functions are sampled independently from a Gaussian process with the squared-exponential kernel (discretized on the uniform 100-point grid) with variance 2 and length scale $0.5$.
-For each parameter function and PDE, the boundary values $u_{0}$ and $u_{1}$ are sampled uniformly in the interval [1,1][-1,1].
-In addition, in the case of the reaction-diffusion PDE, the coefficients $a$ and $c$ are sampled uniformly in the intervals $[0.025,0.075]$ and [2,2][-2,2], respectively.
-Then the PDE is solved for the given parameter function, boundary values, and coefficients to obtain the groundtruth solutions.
-The model was trained by minimizing the mean squared error loss between predicted and groundtruth solutions over a mini-batch of 100 operators sampled randomly from the training data at each iteration, using the Adam optimizer with learning rate 1e-3, for 50k iterations.
-Training took 14min, achieving a training loss 8.4e-3.
-Sample results are shown in Figure 2.
-At inference time, DeepOSets can perform accurate in-context learning, without knowing which PDE, inverse or forward problem, or the value of the equation coefficients and boundary conditions are.
-All these unknowns are correctly disambiguated from the data in the prompt, and the solution corresponding to the query parameter is accurately predicted, without any weight updates.
+To generate the training data for each of the problems, 100 parameter functions are sampled independently from a Gaussian process with the squared-exponential kernel (discretized on the uniform 100-point grid) with variance 2 and length scale $0.5$.
+For each boundary-value problem, the boundary values $u(0)$ and $u(1)$ are sampled uniformly in the interval [1,1][-1,1].
+For the reaction-diffusion problems, the coefficients $a$ and $c$ are sampled uniformly in the intervals $[0.025,0.075]$ and [2,2][-2,2], respectively.
+Then the equation is solved for the given parameters to obtain the ground truth solutions.
+The model was trained by minimizing the mean-squared error loss between predicted and ground truth solutions over a mini-batch of 100 operators sampled randomly from the training data at each iteration, using the Adam optimizer with learning rate 1e-3, for 50k iterations.
+Training the 6-task DeepOSets took only 12min, achieving a training loss 8.4e-3, while the training time of the 12-task DeepOSets was only 38 minutes, achieving a training loss of 4.8e-3, on an Nvidia 4060 Ti GPU with 16GB memory.
+We contrast this to the complex transformer-based architecture in Yang et al.
+((2023b), which requires many hours of training on similar computing hardware.
+Sample results are shown in Figure 2, while MSE values on test data from operators not seen during training are displayed in Table 2.
+DeepOSets achieves accurate predictions without weight updates at inference time, without knowing which task or the values of the equation coefficients and boundary conditions.
+All these unknowns are correctly disambiguated from the examples in the prompt.
+By comparing the 6-task and 12-task errors in Table 2, we observe the interesting fact that the errors are almost universally smaller with the 12-task DeepOSets.
+Considering all 12 tasks simultaneoulsy could be expected to be more challenging due to the larger number of equations but, on the other hand, it presents a larger diversity of solution behaviors.
+Our experiment shows that more diversity of solution operators in the training data is indeed helpful in training a better model that learns to disambiguate the multiple operators more effectively in-context.
+(a) Forward Poisson BVP
+(b) Inverse Poisson BVP
+(c) Forward reaction-diffusion BVP
+(d) Inverse reaction-diffusion BVP
+(e) Forward nonlinear reaction-diffusion BVP
+(f) Inverse nonlinear reaction-diffusion BVP
+Figure 2: Sample results for boundary-value problem test data obtained by the 12-task model.
+In each case, the top plot contains the “condition” (input) functions while the bottom plot contains the corresponding “QoI” (output) functions.
+The gray lines indicate in each case the condition-QoI function pairs given in the prompt.
+The blue line indicates the query parameter function, while the red line indicates the corresponding output of the trained DeepOSets model.
+The black line indicates the groundtruth output corresponding to the query.
+Test Problem 6-task DeepOSets 12-task DeepOSets
+Forward IVP 1 2.65e-03 $\pm$ 9.30e-03
+Inverse IVP 1 1.15e-02 $\pm$ 2.71e-02
+Forward IVP 2 1.65e-03 $\pm$ 8.39e-03
+Inverse IVP 2 1.44e-02 $\pm$ 3.86e-02
+Forward IVP 3 1.19e-02 $\pm$ 6.99e-02
+Inverse IVP 3 1.40e-02 $\pm$ 9.15e-02
+Forward Poisson BVP 3.04e-03 $\pm$ 1.91e-02 4.87e-04 $\pm$ 1.09e-03
+Inverse Poisson BVP 4.60e-03 $\pm$ 1.35e-02 3.84e-03 $\pm$ 1.20e-02
+Forward linear reaction-diffusion BVP 2.11e-03 $\pm$ 4.25e-03 6.49e-04 $\pm$ 9.75e-04
+Inverse linear reaction-diffusion BVP 1.45e-02 $\pm$ 6.01e-02 7.71e-03 $\pm$ 2.75e-02
+Forward nonlinear reaction-diffusion BVP 2.77e-03 $\pm$ 5.69e-03 1.71e-03 $\pm$ 3.32e-03
+Inverse nonlinear reaction-diffusion BVP 1.19e-02 $\pm$ 3.51e-02 1.33e-02 $\pm$ 5.42e-02
+Table 2: Mean Squared Error (MSE) values achieved by the 6-task and 12-task models on test data from each of the 12 tasks individually.
+Next, we evaluated the ability of the DeepOSets framework to extrapolate to sample sizes in inference time that were not seen at training time.
+Indeed, the promised advantages of the DeepOSets approach are its permutation invariance and flexibility regarding the number of examples in the prompt.
+To evaluate the extrapolation capability, we used different number of in-context examples ranging from $N=3$ to $N=20$ at inference time.
+As illustrated in Figure 3, the model successfully generalizes to context sizes significantly larger than those seen in training, and we observe a consistent decrease in MSE as the number of context examples increases.
+This behavior confirms that the model is not overfitting to the specific training sequence length; rather, it effectively performs in-context learning, leveraging additional data points to refine the operator approximation, demonstrating robust generalization.
+(a)   Forward problems
+(b)   Inverse problems
+Figure 3: Zero-shot generalization to larger context sizes.
+The model was trained with a fixed context size of $N=4$.
+During inference, we evaluate performance on context sizes ranging from 3 to 20.
+The results show a clear trend where providing more examples decreases the MSE, indicating the model successfully leverages additional data.
 7 Conclusion
-We have presented DeepOSets for in-context multi-operator learning, a novel non-autoregressive, non-attention-based approach to learning solution operators of multiple parametric PDEs.
+We have presented DeepOSets for in-context multi-operator learning, a novel non-autoregressive, non-attention-based approach to learning solution operators of multiple parametric differential equations.
 By combining DeepSets for set learning with DeepONets for operator learning, DeepOSets addresses key limitations of existing approaches while maintaining strong in-context learning capabilities.
 The proposed architecture offers several advantages: fully parallel processing of prompts enables faster training and inference compared to sequential methods; linear complexity in prompt length provides better scalability than quadratic self-attention mechanisms; and built-in permutation-invariance provides beneficial inductive bias for training efficiency.
-We prove that DeepOSets are universal uniform approximators over a class of operators, which we believe is the first result of its kind to be published in the Scientific Machine Learning area.
-Our experiments on Poisson and reaction-diffusion problems demonstrate that DeepOSets successfully disambiguates between multiple PDEs, forward and inverse problems, and unknown equation coefficients and boundary conditions purely from prompted examples.
-The model achieves accurate predictions without weight updates at inference time, requiring only 14 minutes of training on all four problems simultaneously.
-Future work will extend DeepOSets to higher-dimensional PDEs, more diverse equation families, and longer prompts.
-DeepOSets establishes a computationally efficient foundation for building general-purpose PDE solvers capable of in-context adaptation, contributing toward foundation models for scientific computing.
-Appendix
-Let us begin by proving Lemma 4.1 showing that $(\delta,C)$-discretizations always exist.
+We prove that DeepOSets are universal uniform approximators over a class of operators, which we believe is the first result of its kind to be published in the field of multi-operator learning.
+Our experiments on different tasks from initial-value and boundary-value problems demonstrate that DeepOSets successfully uses the prompted examples to disambiguate between multiple operators corresponding to forward and inverse problems and unknown equation coefficients and boundary conditions.
+The experiments also showed that using a more diverse training set with more solution operators was beneficial to accuracy of in-context learning with the DeepOSets model.
+DeepOSets establishes a computationally efficient foundation for building general-purpose differential equation solvers capable of in-context adaptation, contributing towards establishing foundation models for scientific computing.
+Appendix A Proofs of Theorem 5.1 and Lemma 3.2
+Let us begin by proving Lemma 3.2 showing that $(\delta,C)$-discretizations always exist for any compact metric space $V$.
 Proof of Lemma 3.2.
-Let $K\subset X$ be compact, and let $\delta>0$ be given.
-Choose a maximal $\delta$-separated subset of $K$, i.e., a subset $\mathcal{X}_{\delta}$ such that $d(x_{i},x_{j})\geq\delta$ for all $i\neq j$, and which is maximal with respect to this property.
-It is clear that the balls of radius $\delta$ around the points $x_{i}$ give a minimal covering of $K$, and so by compactness $\mathcal{X}_{\delta}$ must be a finite set.
+Let $V\subset X$ be compact, and let $\delta>0$ be given.
+Choose a maximal $\delta$-separated subset of $V$, i.e., a subset $\mathcal{X}_{\delta}$ such that $d(u_{i},u_{j})\geq\delta$ for all $i\neq j$, and which is maximal with respect to this property.
+It is clear that the balls of radius $\delta$ around the points $u_{i}$ give a minimal covering of $V$, and so by compactness $\mathcal{X}_{\delta}$ must be a finite set.
 We claim that $\mathcal{X}_{\delta}$ is a $(\delta,C)$-covering for any $C>1$.
 First, note that any ball of radius $\delta/2$ contains at most one point of $\mathcal{X}_{\delta}$.
-Indeed, fix $z\in K$ and suppose $x_{i},x_{j}\in\mathcal{X}_{\delta}$ both lie in $B(z,\delta/2)$ with $i\neq j$.
+Indeed, fix $z\in V$ and suppose $u_{i},u_{j}\in\mathcal{X}_{\delta}$ both lie in $B(z,\delta/2)$ with $i\neq j$.
 Then, by the triangle inequality,
-$d(x_{i},x_{j})\;\leq\;d(x_{i},z)+d(z,x_{j})\;<\;\frac{\delta}{2}+\frac{\delta}{2}\;=\;\delta,$
-which contradicts the $\delta$–separation of $\mathcal{X}_{\delta}$. Hence, for every $z\in K$,
-$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(z,\delta/2)\}\bigr|\leq 1.$
-Next, since balls of radius $\delta$ around the points in $\mathcal{X}_{\delta}$ cover $K$, for every $y\in K$ there exists some $x_{j}\in\mathcal{X}_{\delta}$ with $d(y,x_{j})<\delta$, i.e.
-$x_{j}\in B(y,\delta)$.
-Thus, for all $y\in K$,
-$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|\geq 1.$
-Combining these two observations, we obtain, for all $x,y\in K$,
-$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(x,\delta/2)\}\bigr|\;<\;C\,\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|,$
-for any $C>1$. Hence $\mathcal{X}_{\delta}$ is a $(\delta,C)$–discretization of $K$. ∎
-Next, we state a Lemma which we will use in the proof of Theorem 4.2.
-Lemma 6.1.
+$d(u_{i},u_{j})\;\leq\;d(u_{i},z)+d(z,u_{j})\;<\;\frac{\delta}{2}+\frac{\delta}{2}\;=\;\delta,$
+which contradicts the $\delta$–separation of $\mathcal{X}_{\delta}$. Hence, for every $z\in V$,
+$\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(z,\delta/2)\}\bigr|\leq 1.$
+Next, since balls of radius $\delta$ around the points in $\mathcal{X}_{\delta}$ cover $V$, for every $v\in V$ there exists some $u_{j}\in\mathcal{X}_{\delta}$ with $d(v,u_{j})<\delta$, i.e.
+$u_{j}\in B(v,\delta)$.
+Thus, for all $v\in V$,
+$\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(v,\delta)\}\bigr|\geq 1.$
+Combining these two observations, we obtain, for all $u,v\in V$,
+$\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(u,\delta/2)\}\bigr|\;<\;C\,\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(v,\delta)\}\bigr|,$
+for any $C>1$. Hence $\mathcal{X}_{\delta}$ is a $(\delta,C)$–discretization of $V$. ∎
+Next, we state a partition of unity approximation Lemma for a compact subset of $C(K)$, which we will use in the proof of Theorem 5.1.
+Lemma A.1.
 Let $K$ be a compact metric space and $V\subset C(K)$ be compact.
@@
 Proof.
-(of Lemma 6.1) Since $V$ is compact, by the Arzela-Ascoli theorem it is also equicontinuous, and so there exists $\delta>0$ such that if $|x-y|<\delta$, then $|f(x)-f(y)|<\varepsilon$ for all $f\in V$.
+(of Lemma A.1) Since $V$ is compact, by the Arzela-Ascoli theorem it is also equicontinuous, and so there exists $\delta>0$ such that if $|x-y|<\delta$, then $|f(x)-f(y)|<\varepsilon$ for all $f\in V$.
 Since $K$ is also compact, we can pick $y_{1},\cdots,y_{n}\in K$ such that $\{B(y_{i},\delta)\}_{i=1}^{n}$ form a covering of $K$.
@@
 $\begin{split}\|f-g\|_{C(K)}&\leq\|f-\sum_{i=1}^{n}f(y_{i})\varphi_{i}\|_{C(K)}+\|f-g\|_{\ell^{\infty}(\mathbf{y})}\max_{i}\|\varphi_{i}\|_{C(K)}+\|g-\sum_{i=1}^{n}g(y_{i})\varphi_{i}\|_{C(K)}\\ &\leq\|f-g\|_{\ell^{\infty}(\mathbf{y})}+2\varepsilon.\end{split}$ (6)
-We can now return to the proof of Thm 4.2.
+We can now return to the proof of Thm 5.1.
 Proof.
@@
 $\{G(u):~G\in\mathcal{K},~u\in V\}\subset C(K_{2})$ (7)
-is compact. Since $K_{2}$ is also compact, we can apply Lemma 6.1 to get that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
+is compact. Since $K_{2}$ is also compact, we can apply Lemma A.1 to get that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
 $\|G(u_{q})-\sum_{i=1}^{n}G(u_{q})(y_{i})\gamma_{i}\|_{C(K_{2})}<\varepsilon/2,$
@@
 Since $\mathcal{K}$ is compact, it is equi-continuous, so there exists a $\delta$ such that if $\|u-v\|_{C(K_{1})}<25\delta/6$, then $\|G(u)-G(v)\|_{C(K_{2})}<\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$ for every $G\in\mathcal{K}$.
-By Lemma 6.1, there exists $\textbf{x}=\{x_{1},\cdots,x_{k}\}\in K_{1}$ such that
+By Lemma A.1, there exists $\textbf{x}=\{x_{1},\cdots,x_{k}\}\in K_{1}$ such that
 $\|u-v\|_{C(K_{1})}\leq\|u-v\|_{\ell^{\infty}(\mathbf{x})}+\frac{\delta}{6}.$ (10)
@@
 $\Phi(\mathbf{u},\mathbf{v})=\{\varphi_{j}(\mathbf{u})\cdot\mathbf{v},\varphi_{j}(\mathbf{u})\}_{j=1}^{l}.$
-We then have that for any $G\in\mathcal{K}$,
+where $\mathbf{u}\in\mathbb{R}^{k}$ and $\mathbf{v}\in\mathbb{R}^{n}$. We then have that for any $G\in\mathcal{K}$,
 $\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))=\left\{\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\cdot G(u_{i})(\textbf{y}),\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\right\}_{j=1}^{l}$
@@
 which combining with (8) and using (18) gives us,
-$\|G(u_{q})-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
-$\leq\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K)}+\|G(u_{q})(\textbf{y})\cdot\Psi-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
+$\|G(u_{q})-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K_{2})}$
+$\leq\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K_{2})}+\|G(u_{q})(\textbf{y})\cdot\Psi-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K_{2})}$
 $<\varepsilon/2+\varepsilon/2=\varepsilon.$
-Finally, we need to show that $\rho$ and $\Phi$ can be chosen continuous.
+Since $\|.\|_{C(K_{2})}:=\sup_{x\in K_{2}}\|.(x)\|$, this result then holds for any $x_{q}\in K_{2}$, as desired. This completes the main inequality.
+However, we still need to show that $\rho$ and $\Phi$ can be chosen continuous.
 It is clear from the definition that $\Phi$ is continuous.